OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 4 Answer No Solution 1 0 Any Real Number

OpenStudy (amistre64):

divide it all by 4 and see what we get as a result

OpenStudy (anonymous):

I think any real number

OpenStudy (anonymous):

any real number

OpenStudy (lilg132):

how?

OpenStudy (anonymous):

put any real number in the expression, it holds all time..

OpenStudy (anonymous):

infinite number of solutions are there..

OpenStudy (anonymous):

Any real number. Since "x" gets eliminated in the equation, it shows that the eq. is independent of "x". Therefore, no matter what value of "x" you put, you will always get the same answer ( ie. 0=0 ) \[4\left( x+1 \right)=4x+4\] \[4x+4=4x+4\] \[0=0\]

OpenStudy (lilg132):

hmm what does it mean when you have no soloution could you give an example?

OpenStudy (anonymous):

like: 4(x+1)=4x+3

OpenStudy (anonymous):

no number satisfies that

OpenStudy (lilg132):

ok thanks

OpenStudy (anonymous):

Here, Let the problem be : \[x ^{2}=-1\] Find the possible values of "x". Soln : We know that x^2 is always positive, therefore it can never be equal to "-1" (ie. a negative number). Therefore, for this equation, no value of ""x" satisfies it, so no solution.